Math 122 (Section 2)  University of South Carolina
D. Meade  Fall 1997

Quiz 1 : Background Information – Math 122 (Fall 1997)

• Personal Information

Name

Group

Phone Number

E-mail Address

Major

Year

• Mathematical Background

Yes No

Have you taken the Math Placement Exam? Was your placement code B22 or A41? If not, have you received a “C” or better in Math 111 or 115? Do you have prior experience with a graphing calculator? (If so, which one?) Did you (or will you) purchase a calculator for use this semester?

• Other Information

1. Do you have any specific goals or objectives that you hope to fulfill by taking this course?

2. What concerns, if any, do you have about this course?

3. What are your career goals?

4. What are your favorite leisure time activities?